OpenStudy (anonymous):

How much heat energy is required to raise the temperature of 0.367kg of copper from 23.0 ∘C to 60.0 ∘C? The specific heat of copper is 0.0920 cal/(g⋅∘C) .

OpenStudy (aaronq):

use the formula: \(q=m*C_p*\Delta T=m*C_p*(T_f-T_i)\)